OpenStudy (anonymous):

AB and CD are two equal chords of a circle with centre O.OP and OQ are perpendiculars on chords AB and CD respectively.If

OpenStudy (anonymous):

I NEED D FULL PROOF

OpenStudy (anonymous):

Why you need proof?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

75 degrees

OpenStudy (anonymous):

angle APO is 90 degrees OP=OQ as they are perpendiculars to equal chords Hence, in triangle OPQ, angle OPQ = angle OQP Hence OPQ = OQP =15 degrees APQ = APO - OPQ = 90 -15 =75

OpenStudy (anonymous):

thxxxxxxxx